# Efficient Pyramid Channel Attention Network for Pathological Myopia Recognition with Pretraining-and-Finetuning

### v2

@@ 5.6 Limitations
This paper aims to utilize global-local pathology distribution prior to improve the PM recognition performance based on limited available images, which is often ignored by existing deep neural networks in PM recognition tasks. Specifically, we propose an EPCA-Net to recognize PM by adaptively adjusting the relative importance of different scale context features. The results show its great potential to be deployed on ophthalmic equipment, which can be used for clinical PM screening and diagnosis. Moreover, we also found that transferring the pre-trained natural image models to the PM recognition task by medical vision adapter design based on pretraining-and-finetuning paradigm, which provides a novel aspect to enhance PM recognition performance under both limited medical images and limited computational resources. However, there are still some limitations in this paper, as follows:
@@
    We transfer the pre-trained natural image model to the PM recognition task by pretraining-and-finetuning paradigm through adapter design, but pre-trained natural image model used in this paper is pre-trained ResNet50. Combining our EPCA adapter with recent power pre-trained natural image models has not been explored, which may present a useful approach to improve PM recognition performance and other medical image analysis tasks based on limited medical data regime and computational resources.
@@ 6 Conclusion and Future work
In this paper, we propose an EPCA-Net for automatic PM recognition based on fundus images, which adaptively recalibrate the feature maps by mining the global-local pathology distribution information. Given that most existing PM recognition datasets are private, we construct a PM recognition benchmark by collecting fundus images of PM from publicly available datasets, which will be released. The comprehensive experiments on four datasets demonstrate the superiority of our EPCA-Net over state-of-the-art deep neural networks. The EPCA-Net has the potential to be deployed on ophthalmic equipment, which can be a computer-aided diagnosis (CAD) tool for PM screening and diagnosis, advancing precision diagnosis and scientific research development. Moreover, we are the first to adapt the pre-trained natural image models to the PM recognition task by medical vision adapter design with the pretraining-and-finetuning paradigm. This provides a new insight into utilizing large pre-trained natural image models to tackle PM recognition tasks under limited medical images. We hope that our method may prompt the development of domain-knowledge driven-based deep neural network design and adapter design in the PM recognition task and other learning tasks in the limited data regime. The code of this paper will be available at https://github.com/TommyLitlle/EPCANet.
Acknowledgements
This work was supported in part by National Natural Science Foundation of China (No.82272086), Guangdong Provincial Key Laboratory (No.2020B121201001), and Shenzhen Natural Science Fund (No. JCYJ20200109140820699) and the Stable Support Plan Program (No. 20200925174052004)

### v3

@@ 5.6 Limitations
This paper aims to utilize global-local pathology distribution prior to improve the PM recognition performance based on limited available images, which is often ignored by existing deep neural networks in PM recognition tasks. Specifically, we propose an EPCA-Net to recognize PM by adaptively adjusting the relative importance of different scale context features. The results show its great potential to be deployed on ophthalmic equipment, which can be used for clinical PM screening and diagnosis. Moreover, we also find that transferring the pre-trained natural image models to the PM recognition task with medical vision adapter design based on the pretraining-and-finetuning paradigm. This provides a novel aspect to enhance PM recognition performance under both limited medical images and computational resources. However, there are still some limitations in this paper, as follows:
@@
    We transfer the pre-trained natural image model to the PM recognition task by pretraining-and-finetuning paradigm through adapter design, but the pre-trained natural image model used in this paper is pre-trained ResNet50. Combining our EPCA adapter with recent powerful pre-trained natural image models (e.g., vision foundation models) has not been explored, which may present a useful approach to improve PM recognition performance and address other medical image analysis tasks based on limited medical data regime and computational resources.
@@ 6 Conclusion and Future work
In this paper, we propose an EPCA-Net for automatic PM recognition based on fundus images, which adaptively recalibrates the feature maps by mining the global-local pathology distribution information. Considering that most existing PM recognition datasets are private, we construct a PM recognition benchmark by collecting fundus images of PM from publicly available datasets. The comprehensive experiments on four datasets demonstrate the superiority of our EPCA-Net over state-of-the-art deep neural networks. The EPCA-Net has the potential to be deployed on ophthalmic equipment, which has the potential to be a computer-aided diagnosis (CAD) tool for PM screening and diagnosis, advancing precision diagnosis and scientific research development. Moreover, we are the first to adapt the pre-trained natural image models to the PM recognition task by medical vision adapter design with the help of the pretraining-and-finetuning paradigm. This provides new insights into utilizing large pre-trained natural image models to tackle PM recognition tasks under limited medical images. We hope that our method may prompt the development of domain-knowledge driven-based deep neural network design and adapter design in the PM recognition task and other learning tasks in the limited data regime. The code of this paper will be available at https://github.com/TommyLitlle/EPCANet.